OpenStudy (igster1112):

Kate's 85 on her English test was 37 points less than twice the grade on her science test. Is the answer 61 because you do the equation 2x-37=85?

OpenStudy (anonymous):

the equation seems right solving for x (if you are looking for the science grade) 2x-37=85 2x=85+37 x=85+37/2 x=61 looks like you are right (:

OpenStudy (igster1112):

Yay! Thank you 😊

OpenStudy (anonymous):

you're welcome :) nice smiley